# TrajAgent: An LLM-based Agent Framework for Automated Trajectory Modeling via Collaboration of Large and Small Models

### v3

@@ 2.4. Collaborative Learning Schema
2.4.1. Agent Learning via Reasoning
Following Reflexion (Shinn et al., 2024), we design expert optim agent for learning from the experimental records via reasoning. The standard optim agent utilizes the history operation and related results as the feedback to update its action in the next step. Specifically, it works as two stages, including ”think” and ”action”. To support the ”think then action”, it builds a long-term memory for recording all experimental data and a short-term memory for historical actions. In the ”think” stage, optim agent analyzes the long-term memory and meta information of experiment and generating the guidance of action in the short-term memory. In the ”action” stage, optim agent analyze the results in short-term memory to generate the action. Different optimization mechanism utilize the same optim agent with different action space and optimization tips.
Besides, as shown in Figure 1, our framework can also utilize LLM-based agent as the specialized model to complete the trajectory modelling tasks. Thus, we design prompt optimization agent to optimize agent based specialized models. We keep all experimental records in each experiment, including the raw input trajectories, the LLM’s inference results, and the reasoning process. We select the two best-performing trajectories along with their corresponding inference results and reasoning process as few-shot examples, which are then added to the original prompt for a re-run of the experiment. This process of conducting experiments and selecting results is referred to as one iteration. In each iteration, the few-shot examples are updated based on the experiment results.
2.4.2. Model Learning via Training
@@
Data Augmentation. Based on the high-level optim agent, we introduce the specific action space for low-level trajectory data augmentation. Following practice from existing works (Zhou et al., 2024; Dang et al., 2023; Yin et al., 2024), we define a fixed set with ten operators for trajectory data augmentation, e.g., insert, replace, split and so on. During the optimization, the operator set with simple description is provided to the optim agent, it can select any combination of operators with their simple parameters as the action. Then the specialized models are trained with the augmented trajectory data and report performance metrics. The optim agent obtain the feedback information, e.g., performance metrics, from UniEnv to continue update its memory and action.
Parameter Optimization. The action space of parameter optimization is defined based on the parameters of model itself. We define a parameter configuration file for each model, the optim agent reads the configuration file and generates code as the action to update the parameters in it. To better understand the meaning of each parameter, we add comments for each parameter in the file. While the action space is large, this kind of action space is flexible to adapt with any models.
@@
3. Experiments
Table 1. Performance of representative methods across five fundamental trajectory modeling tasks. For the nine subtasks, only one model is presented for each. ‘DA’ represents data augmentation, ‘PO’ denotes parameter optimization, ‘PRO’ denotes prompt optimization, ‘JO’ indicates joint optimization, $\delta$ represents performance improvements.
Task Trajectory Prediction Trajectory Recovery Trajecory Classification Trajectory Trajectory
@@ 3. Experiments
Metric Acc@5 MAE MAE $Acc_{m}$ Acc@5 $Acc_{i}$ AUC MAE Acc@5
Dataset FSQ Porto Porto Tencent FSQ Beijing Porto Earthquake FSQ
Models GETNext LLM-ZS MulT-TTE TrajBERT GraphMM S2TUL LIMP GM-VSAE DSTPP CACSR
Origin 0.3720 0.3110 163.12 42.71 0.2014 0.5755 0.745 0.9892 0.4611 0.31
+DA 0.3894 0.6846 0.3369
+PO 0.3995 128.57 27.78 0.2427 0.757 0.9899 0.3584 0.3466
@@
+JO 0.4002 0.3225 128.57 27.78 0.2427 0.7802 0.7627 0.9899 0.3584 0.3472
$\delta$ 7.58% 3.69% 21.18% 34.96% 20.71% 35.57% 2.38% 0 22.27% 12%
3.1. Settings
Here, we introduce the experimental settings, including the datasets, task models, and evaluation metrics. More detailed information can refer to appendix A.
Data. We utilize the widely used Foursquare (FSQ) (Yang et al., 2016) and Brightkite (BGK) (Cho et al., 2011) in our framework as the default check-in trajectory data. Porto (Kaggle, 2015) and Chengdu (Chuxing, 2018) are integrated in the framework as the default GPS trajectory data. Besides, we use Tencent (Liu et al., 2023) as the road network based methods to support road network based tasks and Beijing (Li et al., 2024) with human labeled intention to support the mobility intention prediction task. Finally, to verify the effectiveness of the whole system, we utilize self-instruct method (Wang et al., 2022) with 5 seed queries to generate 300 user queries as the experiment input.
Models. As shown in Figure 3, our framework supports 18 models spanning 5 core trajectory modeling tasks, which are further categorized into 9 subtasks. The next location prediction task includes FPMC (Rendle et al., 2010), RNN, DeepMove (Feng et al., 2018), GETNext (Yang et al., 2022), LLM-ZS (Beneduce et al., 2024). The travel time estimation task comprises DeepTTE (Wang et al., 2018) and MulT-TTE (Liu et al., 2020a). The trajectory recovery task features TrajBERT (Si et al., 2023), while the map-matching task incorporates DeepMM (Feng et al., 2020) and GraphMM (Liu et al., 2023). The trajectory user linking task includes DPLink (Feng et al., 2019), MainTUL (Chen et al., 2022), and S2TUL (Deng et al., 2023). The mobility intention prediction task is supported by LIMP (Li et al., 2024). The trajectory anomaly detection task employs GM-VSAE (Liu et al., 2020b), and the trajectory generation task includes ActSTD (Yuan et al., 2022) and DSTPP (Yuan et al., 2023). Finally, the trajectory representation method is implemented using CASCR (Gong et al., 2023).
Metrics. We adopt standard practices for each task to select appropriate metrics for evaluating the performance of our framework. The widely used Acc@k metric is employed for next location prediction, map-matching, trajectory user linking, mobility intention prediction, and trajectory representation tasks. The MAE metric is utilized for travel time estimation, trajectory recovery, and trajectory generation tasks. Lastly, the AUC metric is specifically defined for the trajectory anomaly detection task.
3.2. Overall Performance and Generalization Capability of TrajAgent
@@
As shown in Table 1, we select at least one representative models for 9 trajectory modelling tasks to present the effectiveness of proposed framework. As Table 1 shows, TrajAgent supports a variety of widely-known trajectory models and demonstrates superior performance across multiple trajectory modeling tasks and trajectory datasets. The output of TrajAgent consistently outperforms the original methods, achieving performance gains ranging from 2.28% to 35.57%. For instance, in the next-location prediction task, TrajAgent harnesses its agentic workflow and collaborative learning schema for automatic modeling and optimization, leading to significant performance improvements of various models. Specifically, it enhances the deep learning-based model GETNext by 7.58% and the LLM-based model LLM-ZS by 3.69%. Performance gain in other tasks and models are much larger, for example improvement from 20.71% to 34.96% for trajectory recovery tasks. We observe that the improvement in the trajectory anomaly detection task is minimal, primarily due to the strong performance of the original models on the dataset.
As shown in Table 2, we presents a performance comparison for check-in trajectory tasks, including the next-location prediction task and the trajectory user linking task, across three models and two datasets. The first key observation is the consistent improvement observed across tasks, models, and datasets, which highlights the potential generalization of the proposed framework. For different trajectory tasks, datasets, and models, TrajAgent consistently provides transferable performance improvements, ranging from 6.08% to 35.57%. Additionally, we observe that the performance gap between different models (e.g., the top two models for each task) on specific tasks and datasets significantly narrows from 24.88% to 9.5% following the automatic optimization of TrajAgent, emphasizing the critical role of effective data augmentation and parameter optimization. In other words, with the support of TrajAgent, we can easily obtain more robust and reliable task performance with less relying on the delicate model designs, data processing and parameter selection.
Table 2. Comparison of task performance on check-in trajectories for TrajAgent with different configurations, which demonstrate the generalization of TrajAgent across different tasks, models and datasets. ‘DA’ represents Data Augmentation, ‘PO’ denotes Parameter Optimization, ‘JO’ indicates Joint Optimization. $\delta$ represents improvements.
Task Next Location Prediction Trajectory User Linking
Model RNN DeepMove GETNext MainTUL DPLink S2TUL
Metrics Acc@5 Hit@5
FSQ Origin 0.1795 0.3422 0.3720 0.4871 0.7551 0.5755
+DA 0.2667 0.4018 0.3894 0.5973 0.7551 0.6846
+PO 0.1795 0.3422 0.3995 0.5691 0.7686 0.7570
+JO 0.2717 0.4018 0.4002 0.6121 0.8010 0.7802
$\delta$ 51.36% 17.42% 7.58% 25.66% 6.08% 35.57%
BGK Origin 0.4422 0.5570 0.5324 0.5908 0.8993 0.5802
+DA 0.5416 0.5647 0.6026 0.6836 0.9613 0.6903
+PO 0.5022 0.6041 0.6116 0.6683 0.9552 0.7137
+JO 0.5470 0.6100 0.6227 0.7145 0.9622 0.7240
$\delta$ 23.70% 9.52% 16.96% 20.94% 6.99% 24.78%
Table 3. Execution success rates and task performance at each stage of the agentic workflow of TrajAgent across different LLMs. The Acc@5 is obtained by evaluating the next-location prediction task, with DeepMove as the default specialized model.
@@
Table 4. Ablation study of TrajAgent. ‘MS’ stands for Model Selection, ‘DA’ represents Data Augmentation, ‘PO’ denotes Parameter Optimization, ‘JO’ stands for Joint Optimization. $\downarrow$ indicates a decrease in performance, and $\uparrow$ indicates an improvement.
@@
w/o Reflection 100% 95$\downarrow$ 98% 0.3028$\downarrow$ 90$\uparrow$ 0.1872$\downarrow$ 82$\downarrow$ 0.3212$\downarrow$
w/o Memory 99$\downarrow$ 80$\downarrow$ 85$\downarrow$ 0.1707$\downarrow$ 70$\downarrow$ 0.2050$\uparrow$ 68$\downarrow$ 0.1804$\downarrow$
3.3. Ablation Study and Parameter Analysis of the Agentic Workflow
@@
Table 3 compares the execution efficiency of TrajAgent implemented by different LLMs, across each stages in the trajectory modelling workflow. We can observe that Qwen2-72B and GPT-4o-mini demonstrate the highest success rates (i.e., Succ.) across key stages such as Data Extraction, Processing, and Data/Model Selection, with over 90% success in each. They also excel in Joint Optimization, with success rates of 94% and 96%, resulting in superior Acc@5 scores of 0.4333 and 0.3724 respectively. In contrast, models like Gemma-2-9B and LLama3-8B struggle with lower processing and data selection success rates, which results in reduced overall performance. Their weaker performance in key optimization stages, especially in parameter selection, reflects their limitations in effectively supporting TrajAgent. These results show that TrajAgent’s efficiency is strongly influenced by the base LLM, with high-performing models like Qwen-72B and GPT-4o-mini significantly enhancing its capabilities. To fully realize TrajAgent’s potential, it is essential to improve the LLM performance in data augmentation and parameter optimization, as many LLMs currently struggle in these aspects, which is the important future direction of TrajAgent.
@@
To assess the effectiveness of the key components in TrajAgent, we conducted an ablation study using execution efficiency as the evaluation criterion. Specifically, we examined the framework by isolating two main designs: the memory unit and the reflection mechanism, resulting in two variants: 1) w/o Reflection, where the reflection mechanism is removed, and 2) w/o Memory, where the memory unit is excluded. Additionally, it is important to note that GPT-4o-mini was used as the base LLM for all variants of TrajAgent in this study. The experimental results are presented in Table 4. We can find that: 1) Removing either component leads to average performance declines, with the memory unit being especially critical for maintaining execution efficiency. For example, removing the memory unit causes significant drops in accuracy for both Data Augmentation (DA) and Parameter Optimization (PO), underscoring its essential role in these processes. 2) Interestingly, removing a component occasionally results in slight increases in success or accuracy, suggesting that some components may introduce overhead or complexity in specific stages. Overall, the combined use of the memory and reflection mechanisms is crucial for optimizing TrajAgent’s performance.
Due to the importance of reflection and memory in the TrajAgent, we analyze the effects of two important related parameters: 1) thought step: the number of steps that agent thought before taking action in reflection, and 2) memory size: the size of memory units in TrajAgent. As shown in Figure LABEL:fig:param-test:thought, increasing thought steps significantly enhances Joint Optimization (JO) and Data Augmentation (DA). This indicates that giving TrajAgent more reasoning time improves its overall performance. However, the peak performance for Parameter Optimization (PO) and JO at around 10 steps suggests that once sufficient contextual understanding is achieved, further increasing the thought steps provides diminishing returns and may not be beneficial. From Figure LABEL:fig:param-test:memory, we find that increasing memory size initially improves performance for JO and DA, with peak accuracy reached at a memory size of 10. This suggests that having sufficient memory allows TrajAgent to store and utilize relevant information effectively. However, beyond this point, performance starts to decline, indicating that excessive memory may introduce unnecessary complexity or irrelevant data, leading to diminished returns for JO and DA. For Parameter Optimization (PO), memory size has little effect, reflecting that this task does not rely as heavily on memory records.
3.4. Comparison with Automated Methods
Figure 5. Compared to Optuna and UrbanLLM, TrajAgent achieves better performance in trajectory user linking tasks with S2TUL as the specialized model. Here, POx represents parameter optimization performed over x steps, while DAx refers to data augmentation conducted over x steps.

### v5

@@ 2.4 Collaborative Learning Schema
2.4.1 Agent Learning via Reasoning
Following Reflexion shinn2024reflexion , we design expert optim agent for learning from the experimental records via reasoning. The standard optim agent utilizes the history operation and related results as the feedback to update its action in the next step. Specifically, it works as two stages, including "think" and "action". To support the "think then action", it builds a long-term memory for recording all experimental data and a short-term memory for historical actions. In the "think" stage, optim agent analyzes the long-term memory and meta information of experiment and generating the guidance of action in the short-term memory. In the "action" stage, optim agent analyze the results in short-term memory to generate the action. Different optimization mechanism utilize the same optim agent with different action space and optimization tips. Besides, as shown in Figure 1, our framework can also utilize LLM-based agent as the specialized model to complete the trajectory modelling tasks. Thus, we design prompt optimization agent to optimize agent based specialized models. We keep all experimental records in each experiment, including the raw input trajectories, the LLM’s inference results, and the reasoning process. We select the two best-performing trajectories along with their corresponding inference results and reasoning process as high-scoring memory entries, which are then added to the original prompt for a re-run of the experiment. This process of conducting experiments and selecting results is referred to as one iteration. In each iteration, the high-scoring memory entries are updated based on the experiment results.
[…]
2.4.2 Model Learning via Training
@@
Data Augmentation. Based on the high-level optim agent, we introduce the specific action space for low-level trajectory data augmentation. For GPS/map-based trajectories, we adopt a geometry-aware augmentation pipeline inspired by DeepMM [12]: (1) raw trajectories are first downsampled to preserve spatial topology; (2) noise is injected only within road network constraints; (3) augmented samples are validated for temporal consistency before merging with original data. For check-in trajectories, we follow the practice from existing works zhou2024contrastive ; dang2023uniform ; yin2024dataset , defining a fixed set with ten operators for trajectory data augmentation, e.g., insert, replace, split and so on. During the optimization, the operator set with simple description is provided to the optim agent, it can select optimal operator sequence(combination of operators with their simple parameters) and parameter configuration as the action, guided by training feedback. Then the specialized models are trained with the augmented trajectory data and report performance metrics. The optim agent obtain the feedback information, e.g., performance metrics, from UniEnv to continue update its memory and action.
Parameter Optimization. The action space of parameter optimization is defined based on the parameters of model itself. We define a parameter configuration file for each model, the optim agent reads the configuration file and generates code as the action to update the parameters in it. To better understand the meaning of each parameter, we add comments for each parameter in the file. This kind of action space is flexible to adapt with any models.
@@
3 Experiments
3.1 Settings
Data. We utilize the widely used Foursquare (FSQ) yang2016participatory and Brightkite (BGK) cho2011friendship in our framework as the default check-in trajectory data. Porto kaggle2015taxitrajectory and Chengdu didi2018gaia are integrated in the framework as the default GPS trajectory data. Besides, we use Tencent liu2023graphmm as the road network based methods to support road network based tasks and Beijing li2024limp with human labeled intention to support the mobility intention prediction task. Finally, to verify the effectiveness of the whole system, we utilize self-instruct method wang2022self with 5 seed queries to generate 300 user queries as the experiment input.
Models. As shown in Figure 3, our framework supports 18 models spanning 5 core trajectory modeling tasks, which are further categorized into 9 subtasks. The next location prediction task includes FPMC rendle2010factorizing , RNN, DeepMove feng2018deepmove , GETNext yang2022getnext , LLM-ZS beneduce2024large . The travel time estimation task comprises DeepTTE wang2018will and MulT-TTE liu2020multi . The trajectory recovery task features TrajBERT si2023trajbert , while the map-matching task incorporates DeepMM feng2020deepmm and GraphMM liu2023graphmm . The trajectory user linking task includes DPLink feng2019dplink , MainTUL chen2022mutual , and S2TUL deng2023s2tul . The mobility intention prediction task is supported by LIMP li2024limp . The trajectory anomaly detection task employs GM-VSAE liu2020online , and the trajectory generation task includes ActSTD yuan2022activity and DSTPP yuan2023spatio . Finally, the trajectory representation method is implemented using CASCR gong2023contrastive .
Metrics. We adopt standard practices for each task to select appropriate metrics for evaluating our framework. The widely used Acc@k metric is employed for next location prediction, map-matching, trajectory user linking, mobility intention prediction, and trajectory representation tasks. The MAE metric is utilized for travel time estimation, trajectory recovery, and trajectory generation tasks. Lastly, the AUC metric is specifically defined for the trajectory anomaly detection task.
Table 1: Performance of representative methods across five fundamental trajectory modeling tasks. For the ten subtasks, only one model is presented for each. ‘DA’ represents data augmentation, ‘PO’ denotes parameter optimization, ‘PRO’ denotes prompt optimization, ‘JO’ indicates joint optimization, $\delta$ represents performance improvements.
Task Trajectory Prediction Trajectory Recovery Trajectory Classification Trajectory Trajectory
@@ 3.1 Settings
Metric Acc@5 Acc@5 MAE MAE MAE $Acc_{m}$ Acc@5 $Acc_{i}$ AUC MAE Acc@5
Dataset FSQ FSQ Porto Tencent Porto Tencent FSQ Beijing Porto Earthquake FSQ
Models GETNext LLM-ZS MulT-TTE DutyTTE TrajBERT GraphMM S2TUL LIMP GM-VSAE DSTPP CACSR
Origin 0.3720 0.3110 163.12 190.82 42.71 0.2014 0.5755 0.745 0.9892 0.4611 0.31
+DA 0.3894 0.3258 0.6846 0.3369
+PO 0.3995 0.3302 128.57 179.01 27.78 0.2427 0.757 0.9899 0.3584 0.3466
@@
+JO 0.4002 0.3350 128.57 179.01 27.78 0.3422 0.7802 0.7627 0.9899 0.3584 0.3472
$\delta$ 7.58% 7.72% 21.18% 6.19% 34.96% 69.91% 35.57% 2.38% 0 22.27% 12%
Table 2: Comparison of performance on check-in trajectories for TrajAgent with different configurations, which demonstrate the generalization of TrajAgent across different tasks, models and datasets.
FSQ BGK
Task Next Location Prediction Trajectory User Linking Next Location Prediction Trajectory User Linking
Model RNN DeepMove GETNext MainTUL DPLink S2TUL RNN DeepMove GETNext MainTUL DPLink S2TUL
Metrics Acc@5 Hit@5 Acc@5 Hit@5
Origin 0.1795 0.3422 0.3720 0.4871 0.7551 0.5755 0.4422 0.5570 0.5324 0.5908 0.8993 0.5802
+DA 0.2667 0.4018 0.3894 0.5973 0.7551 0.6846 0.5416 0.5647 0.6026 0.6836 0.9613 0.6903
+PO 0.1795 0.3422 0.3995 0.5691 0.7686 0.7570 0.5022 0.6041 0.6116 0.6683 0.9552 0.7137
+JO 0.2717 0.4018 0.4002 0.6121 0.8010 0.7802 0.5470 0.6100 0.6227 0.7145 0.9622 0.7240
$\delta$ 51.36% 17.42% 7.58% 25.66% 6.08% 35.57% 23.70% 9.52% 16.96% 20.94% 6.99% 24.78%
3.2 Overall Performance and Generalization Capability of TrajAgent
@@
As shown in Table 1, we select at least one representative models for 9 trajectory modelling tasks to present the effectiveness of proposed framework. As Table 1 shows, TrajAgent supports a variety of widely-known trajectory models and demonstrates superior performance across multiple trajectory modeling tasks and trajectory datasets. The output of TrajAgent consistently outperforms the original methods, achieving performance gains ranging from 2.28% to 69.91%. For instance, in the next-location prediction task, TrajAgent harnesses its agentic workflow and collaborative learning schema for automatic modeling and optimization, leading to significant performance improvements of various models. Specifically, it enhances the deep learning-based model GETNext by 7.58% and the LLM-based model LLM-ZS by 7.72%. Performance gain in other tasks and models are much larger, for example improvement from 34.96% to 69.91% for trajectory recovery tasks. We observe that the improvement in the trajectory anomaly detection task is minimal, primarily due to the strong performance of the original models on the dataset.
As shown in Table 2, we presents a performance comparison for check-in trajectory tasks, including the next-location prediction task and the trajectory user linking task, across three models and two datasets. The first key observation is the consistent improvement observed across tasks, models, and datasets, which highlights the potential generalization of the proposed framework. For different trajectory tasks, datasets, and models, TrajAgent consistently provides transferable performance improvements, ranging from 6.08% to 35.57%. Additionally, we observe that the performance gap between different models (e.g., the top two models for each task) on specific tasks and datasets significantly narrows from 24.88% to 9.5% following the automatic optimization of TrajAgent, emphasizing the critical role of effective data augmentation and parameter optimization.
Table 3: Execution success rates and task performance at each stage of the agentic workflow of TrajAgent across different LLMs. The Acc@5 is obtained by evaluating the next-location prediction task, with DeepMove as the default specialized model.
@@
Table 4: Ablation study of TrajAgent. ‘MS’ stands for Model Selection, ‘DA’ represents Data Augmentation, ‘PO’ denotes Parameter Optimization, ‘JO’ stands for Joint Optimization. $\downarrow$ indicates a decrease in performance, and $\uparrow$ indicates an improvement.
@@
[…]
(a) (a-b) The impact of thought steps and memory size when using DeepMove in next location prediction tasks. (c) Compared to Optuna and UrbanLLM, TrajAgent achieves better performance in trajectory user linking tasks with S2TUL as the specialized model.
3.3 Ablation Study and Parameter Analysis of the Agentic Workflow
@@
Table 3 compares the execution efficiency of TrajAgent implemented by different LLMs, across each stages in the trajectory modelling workflow. We can observe that Qwen2-72B and GPT-4o-mini demonstrate the highest success rates (i.e., Succ.) across key stages such as Data Extraction, Processing, and Data/Model Selection, with over 90% success in each. In contrast, models like Gemma-2-9B and LLama3-8B struggle with lower processing and data selection success rates, which results in reduced overall performance. Their weaker performance in key optimization stages, especially in parameter selection, reflects their limitations in effectively supporting TrajAgent. These results show that TrajAgent’s efficiency is strongly influenced by the base LLM, with high-performing models like Qwen-72B and GPT-4o-mini significantly enhancing its capabilities.
@@
We conducted an ablation study by isolating two main designs: the memory unit and the reflection mechanism, resulting in two variants: 1) w/o Reflection, where the reflection mechanism is removed, and 2) w/o Memory, where the memory unit is excluded. The experimental results are presented in Table 4. We can find that: 1) Removing either component leads to average performance declines, with the memory unit being especially critical for maintaining execution efficiency. 2) Interestingly, removing a component occasionally results in slight increases in success or accuracy, suggesting that some components may introduce overhead or complexity in specific stages. Overall, the combined use of the memory and reflection mechanisms is crucial for optimizing TrajAgent’s performance.
Due to the importance of reflection and memory in the TrajAgent, we analyze the effects of two important related parameters: 1) thought step: the number of steps that agent thought before taking action in reflection, and 2) memory size: the size of memory units in TrajAgent.
As shown in Figure LABEL:fig:param-test:thought, performance initially improves with more thought steps, reflecting enhanced reasoning depth. However, beyond a threshold (e.g., 20 steps), performance declines—likely due to overfitting to suboptimal action sequences or repetitive exploration without sufficient novelty. Similarly, in Figure LABEL:fig:param-test:memory, increasing memory size beyond 10 leads to performance degradation, suggesting that excessive historical records may introduce noise or bias the agent toward previously failed strategies (a phenomenon we term the ‘optimization trap’).
3.4 Analysis of Optimization Failure Modes and Improvements
While TrajAgent demonstrates strong performance across diverse trajectory tasks, we observe that its optimization efficacy does not monotonically improve with increasing reasoning depth or memory capacity. To understand this phenomenon and enhance robustness, we conduct an in-depth analysis of the underlying optimization dynamics.
Optimization Trap in Long Reasoning Chains: As the number of thought steps increases, the agent may converge prematurely to a local optimum. Once trapped, the agent stops exploring novel strategies and repeatedly refines the same ineffective combination. This behavior is exacerbated in weaker LLMs, which lack sufficient reasoning capacity to escape such traps. In contrast, reasoning models like DeepSeek-v3 exhibit more diverse exploration and are more likely to discover globally superior configurations within fewer steps (see Table 10 in Appendix).
Memory Saturation and Noise Accumulation: Similarly, increasing memory size improves exploration efficiency up to a point, but larger memories introduce noisy or redundant historical records. Low-performing action sequences, if retained, can bias future decisions and lead the agent to revisit failed strategies—especially when memory is not actively curated. This “memory pollution” effect explains the performance drop in Figure LABEL:fig:param-test:memory. Crucially, stronger reasoning models (e.g., DeepSeek-v3, Gemini-2.0) are less affected, as shown in Table 10 in Appendix: they maintain or even improve performance with larger memories by better filtering useful experiences, whereas weaker models degrade significantly. This highlights memory content management and optimization as a key factor in mitigating memory pollution.
It is also worth noting that although Figure LABEL:fig:param-test:memory shows parameter optimization (PO) is less sensitive to memory size than data augmentation (DA), excessive memory (>10 entries) still degrades PO due to noise accumulation. The apparent stability in PO stems from its smaller action space, but memory pruning is equally critical for both.
Mitigation via Contrastive Reflection and Memory Pruning: To address these issues, we introduce two practical improvements: (1) Contrastive Reflection: During the reflection phase, the agent explicitly compares successful and failed trials, adjusting operator parameters to avoid repeating ineffective combinations. This encourages diverse yet informed exploration. (2) Dynamic Memory Pruning: We periodically discard low-scoring memory entries and retain only high-performing trajectories as high-scoring memory entries to guide future planning. As shown in Tables 12 and 11 in Appendix, these strategies significantly stabilize optimization.Similarly, memory pruning enables consistent gains across memory lengths, with performance no longer collapsing at large capacities.
3.5 Comparison with Automated Methods